Opened 11 years ago

Closed 10 years ago

#94 closed flight processing (fixed)

IPY07/10, flight day 203/2007, Russell Glacier

Reported by: mggr Owned by: mggr
Priority: alpha 4 high Milestone: 2007 data processing completion
Component: Processing: general Keywords: IPY
Cc: Other processors:

Description (last modified by mggr)

Data location: ~arsf/arsf_data/in_progress/2007/flight_data/ipy/IPY07_10-2007_203_Russell_Glacier

Data arrived from ARSF via SATA disk transfer in August, with post-processed GPS arriving in December following Kidlington visit.

Scientific details: (details of IPY applications not currently available at PML)

PI: Andrew Shepherd

Sensors:

  • ATM

Note there appears to also be Eagle and Hawk data on this flight that isn't mentioned on the logsheet.

Quicklook

TBD.

Change History (9)

comment:1 Changed 11 years ago by mggr

Process Eagle and Hawk to compare it against the ATM - lets see if it is associated or not, then query Gary.

comment:2 Changed 11 years ago by anee

  • Owner changed from mggr to mark1

Processed ATM flight data located in: /users/rsg/anee/scratch/IPY07_10-2007_203_Russell_Glacier

Delivery directory created

comment:3 Changed 11 years ago by mark1

  • Owner changed from mark1 to anee

Has been checked and example commands tested for each sensor, data looks ok. All lev1 files checked and ok. No mosaic images.

rsynced to /users/rsg/arsf/arsf_data/in_progress/2007/flight_data/ipy/IPY07_10-2007_203_Russell_Glacier

rsync mosaics over and then it will be ready to deliver.

comment:4 Changed 11 years ago by mggr

  • Description modified (diff)

Misdelivered due to the wrong PI name on the logsheet - needs redelivery to Andrew Shepherd (http://www.geos.ed.ac.uk/people/person.html?indv=1205)

comment:5 Changed 11 years ago by mggr

Verify overflows on Eagle and Hawk, and reprocess if necessary.

comment:6 Changed 11 years ago by mggr

PI stated he doesn't want a delivery of the data.

Please reprocess the Eagle and Hawk anyway, and prepare a redelivery (just copy the existing delivery and overwrite the E/H data) - we still need to deliver the corrected versions to NEODC.

comment:7 Changed 11 years ago by mggr

Eagle and Hawk did show (minor) overflows in non-cloud areas, reprocessed.

NERC requires us to deliver, so sending on a disk today.

comment:8 Changed 11 years ago by anee

  • Owner changed from anee to mggr

comment:9 Changed 10 years ago by benj

  • Resolution set to fixed
  • Status changed from new to closed
Note: See TracTickets for help on using tickets.